# Geometric Origin of Quantum Entanglement

### v1

@@ Abstract
We investigate massless representations related to the extension of Poincarè group constructed in [1]. These representations differ from Wigner’s ones of standard Poincarè group because the stabilizer of a lightlike momentum in the extended group is $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}Z_{2}$, with factor $Z_{2}=\{1,-1\}$ generated by involution $\Lambda_{\infty}(\theta,\phi)$ which represents infinite velocity limit of a superluminal boost along spatial direction identified by polar and azimuthal angles $\theta,\phi$. The unitary irreducible representations (UIRs) of massless particles in this extension must decompose as a direct sum of a massless forward (positive zeroth component momentum) and massless backward (negative zeroth component momentum) Wigner’s representations linked by internal two valued degree of freedom given by the two possible eigenvalues of $U(\Lambda_{\infty})$. We prove that these representations are unitarily equivalent to entangled states of two qubits. This provides a geometric origin of quantum entanglement for photons in the framework of quantum field theory: photons appear as superpositions of backward and forward propagating electromagnetic waves depending on the eigenvalue of $U(\Lambda_{\infty})$ and this dependency gives rise to correlations between the values of local observables identical to those experienced with an entangled state of two qubits. Finally we describe an experiment capable of distinguishing the two eigenvalues of $U(\Lambda_{\infty})$ providing experimental falsification of the theory.
@@ 1 Introduction
Quantum entanglement is usually introduced as a feature of composite quantum systems: given a tensor product $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, there exist pure states that cannot be written as product vectors. In nonrelativistic quantum mechanics this structure is taken as primitive. By contrast, in relativistic quantum theory the state space is constrained by spacetime symmetry: one starts from unitary (or projective) representations of the Poincaré group and builds quantum fields and particles from them. It is then natural to ask whether some instances of entanglement may have a direct geometric origin in the representation theory of spacetime symmetries.
@@
$\mathcal{L}_{\mathrm{ext}}\;\cong\;\bigl(\mathrm{SO}(3,1)^{+}_{\uparrow}\rtimes_{\mathrm{Ad}_{\Lambda_{\infty}}}Z_{2}\bigr)\times Z_{2}$
shares the same identity component as the ordinary Lorentz group $O(3,1)$ but has disconnected components generated by $\Lambda_{\infty}(\theta,\phi)$ and $\Lambda_{-\infty}(\theta,\phi)$ in place of parity and time-reversal. When translations are included, this leads to an extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{1,3}\rtimes\mathcal{L}_{\mathrm{ext}}$ whose unitary irreducible representations (UIRs) differ from the standard Wigner classification in two essential ways: (i) timelike and spacelike orbits are merged into a single tachyon/massive multiplet, and (ii) in the massless sector the stability subgroup is enlarged from Wigner’s $ISO(2)$ to
$ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}Z_{2},\qquad Z_{2}=\{1,-1\},\quad\Lambda_{-\infty}^{2}=\mathbb{I}.$
The presence of this extra $Z_{2}$ factor in the stabiliser of a lightlike orbit of the extended group has a direct and unavoidable representation–theoretic consequence. Every massless UIR of $\mathcal{P}_{\mathrm{ext}}$ is no longer a single forward Wigner representation, but comes as a two components object
@@
where $\pi^{\text{fwd}}$ and $\pi^{\text{vwd}}$ are forward and backward lightlike Wigner’s UIRs of ordinary Poincaré group, and $\varepsilon$ is the eigenvalue of the unitary operator $U(\Lambda_{\infty})$. From the extended group viewpoint, the massless sector always carries an additional binary internal label $\varepsilon=\pm 1$ attached to the pair $(\pi^{+},\pi^{-})$.
The central observation of the present paper is that this binary structure realises, in a spacetime–geometric way, the algebraic pattern usually associated with entangled qubit pairs. The representation space of $\pi_{\mathrm{ml}}^{\mathrm{ext}}$ can be written as a direct sum
@@
where the above summands carry forward and backward massless Wigner UIRs respectively, while the action of $U(\Lambda_{\infty})$ exchanges the two sectors up to a sign determined by $\varepsilon$. We show that there exist:
@@
The aim of this paper is to make this equivalence precise and to explore its physical implications. We first recall the construction of the extended Lorentz and Poincaré groups and the structure of their massless UIRs. We then make explicit the isometry between the direct–sum representation $\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ and the tensor–product representation $\mathcal{H}\otimes\mathbb{C}^{2}$, and we construct the associated *–homomorphism of local observable algebras. Finally, we outline a local–tomography experiment, within standard quantum optics and interferometry, that should detect the additional binary degree of freedom $\varepsilon$ and thus provide an empirical test of the extended Poincaré symmetry underlying this geometric origin of quantum entanglement.
2 UIRs of Extended Poincaré Group
In this section we briefy recall the mathematical construction leading to the extension of Poincarè group including superluminal observers.
@@ 2 UIRs of Extended Poincaré Group
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. Including translations we defined the extended Poincarè group
@@
on UIRs of Poincarè group $\mathcal{P}_{0}$. In order to explain this we call $N\equiv\mathcal{P}_{0}=(\mathcal{T}\rtimes SO(3,1)^{+}_{\uparrow})$ and define
@@
An element $z\in Z$ acts on $N$ via automorphisms, namely its action on a given element of the Poincarè group $n=(h,a)\in N$ with $h\in SO(3,1)^{+}_{\uparrow}$:
@@
This induces an action on the set of charachters (i.e. UIRs) of $\mathcal{P}_{0}$:
@@
The set $\hat{N}$ is constituted by UIRs of Poincarè group which have been classified by Wigner in [2]. UIRs of Poincarè group are classified by the value of the four momentum vector $p=p_{\mu}$ together with the value of the invariant lenght $|L_{s}\cdot p|^{2}$ associated to the orbit
@@
$p$ is called the representative of the induced UIR of Poincarè group. Given choice of $p_{\mu}$ an $SO(3,1)^{+}_{\uparrow}$-orbit representative in Wigner’s classification, its invariant could be $p^{\mu}p_{\mu}\neq 0$ or $p^{\mu}p_{\mu}=0$. Choosing a non massless ($p^{\mu}p_{\mu}\neq 0$) representative $p_{\mu}$ we have:
@@
2.1 Non Massless Representations
For a given choice of momentum $p_{\mu}$, the possible non-lightlike $z$-orbits of the extended group are those with $p^{\mu}p_{\mu}\neq 0$ and give rise to the following UIRs of the Poincarè group $\mathcal{P}_{0}$:
$\hat{N}=\{\pi^{+}_{\text{mass}},\pi^{-}_{\text{mass}},\pi_{\text{tach}}\}$ (18)
The stabilizers of the corresponding orbits in (15) $Z_{\pi}={e}$, $\forall\pi\in\hat{N}$ and from (17)we have:
$U(\Lambda_{\infty})\pi^{+}_{\text{mass}}(n)U(\Lambda_{\infty})^{-1}=\pi^{+}_{\text{mass}}(\Lambda_{\infty}n\Lambda_{\infty}^{-1})=\pi_{\text{tach}}$ (19)
$U(\Lambda_{-\infty})\pi^{+}_{\text{mass}}(n)U(\Lambda_{-\infty})^{-1}=\pi^{-}_{\text{mass}}(\Lambda_{-\infty}n\Lambda_{-\infty}^{-1})=\pi_{\text{tach}}$ (20)
$U(-\mathbb{I})\pi^{+}_{\text{mass}}(n)U(-\mathbb{I})^{-1}=\pi^{+}_{\text{mass}}((-\mathbb{I})n(-\mathbb{I}))=\pi^{-}_{\text{mass}}$ (21)
We thus have that the set of UIRs in (18) is a set of equivalent irreducible representations of the extended group and in order to specify the full UIR we need to derive the action of the operators $\bar{U}(z)$ for $z$ in $Z$ on wavefunctions. Again the trick is to start from the known Wigner’s representations for $z=I$ and derive the actions of $\bar{U}(z)$ for $z\neq I$ from the action of the extension group $Z$ on Wigner’s representations space. In [1] it is showed:
$\bar{U}(\Lambda_{\infty})\psi](p)=\psi(\Lambda_{\infty}^{-1}p)$ (22)
and similarly for $\bar{U}(-I)$ and $\bar{U}(-\Lambda_{\infty})$we have:
$\bar{U}(-I)\psi](p)=\psi(-p)\;\;\;\;[\bar{U}(-\Lambda_{\infty})\psi](p)=\psi((-\Lambda_{\infty})^{-1}p)$ (23)
2.2 Massless Representations
In [1] it is considered the lightlike zz–orbit
@@ 2.2 Massless Representations
thus interchanging the forward and backward light–cones. In particular $\Lambda_{\infty}$ does not belong to the geometric stabilizer of $p_{0}$. However, at the level of massless UIRs of $\mathcal{P}_{\text{ext}}$, $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent since the unitary operator representing $U(\Lambda_{\infty})$ is thus an intertwiner between $\pi_{\text{fwd}}$ and $\pi_{\text{bwd}}$
@@
Thus, under the action of the discrete factor $Z_{(\theta,\phi)}=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ on the dual space $\widehat{\mathcal{P}_{0}}$, (i.e. the set of Wigner’s UIRs) the set $\Pi\equiv\{\pi^{\text{fwd}},\pi^{\text{bwd}}\}$ constitute a single equivalence class of UIRs in which the transformation $U(\Lambda_{\infty})$ is a non trivial stabilizer, namely:
@@
where $C$ implements the intertwining between $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ as above. One checks, using a section $k_{0}$ adapted as
@@
so $\bar{U}(\Lambda_{\infty})$ correctly represents the extension.
@@
3 Equivalence of Massless UIRs with Entangled Two–Qubits States
@@
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=(I\otimes\sigma_{x})\big(\psi_{\text{fwd}}\otimes|\text{bwd}\rangle+\psi_{\text{bwd}}\otimes|\text{fwd}\rangle\big)$ (45)
@@
Since this must hold for all wavefunctions in $\mathcal{H}_{\oplus}$ we conclude that, via the isometric identification $V$, the sector–swap operator $U(\Lambda_{\infty})$ on $\mathcal{H}\oplus\mathcal{H}$ is unitarily equivalent to the operator $\mathbf{1}\otimes\sigma_{x}$ on $\mathcal{H}\otimes\mathbb{C}^{2}$.
4 Experimental Proposal: Tomography of $\varepsilon$ Eigenvalue
@@ 4.1 Specialisation to a fixed direction and single–mode photons
  • Choice of $\hat{n}$. Fixing $(\theta,\phi)$ amounts to choosing a physical propagation axis in space. In the proposed single–photon experiment we specialise, without loss of generality, to $\hat{n}=(0,0,1)$, i.e. the optical axis is aligned along the zz–direction. The interferometer arms then realise the two counterpropagating lightlike modes with momenta
@@
    which are the forward and backward representatives associated with this choice of $\hat{n}$. Rotating the entire setup in space would simply correspond to choosing a different pair $(\theta,\phi)$ in the abstract construction.
so $p_{0}=\omega(1,0,0,1),$ with $\Lambda_{\infty}=\Lambda_{\infty}(0,0)$ an involution exchanging the time coordinate $t$ with the spatial coordinate $z$. we can represent a fixed helicity component by kets

### v2

# Geometric Quantum Computation

@@ Abstract
Massless unitary irreducible representations of the extension of Poincarè group constructed in [1] differ from Wigner’s ones of standard Poincarè group because the stabilizer of a lightlike momentum in the extended group is $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}\mathbb{Z}_{2}$, with factor $\mathbb{Z}_{2}=\{1,-1\}$ generated by involution $\Lambda_{\infty}(\theta,\phi)$ which represents infinite velocity limit of a superluminal boost along spatial direction identified by polar and azimuthal angles $\theta,\phi$. These must decompose as a direct sum of a massless forward (positive zeroth component momentum) and massless backward (negative zeroth component momentum) Wigner’s representations linked by internal two valued degree of freedom given by the two possible eigenvalues of $U(\Lambda_{\infty})$. We prove that these representations are operationally equivalent (on local observables) to entangled states of a two-qubit system. This provides a geometric origin of quantum entanglement for photons in the framework of quantum field theory: photons appear as superpositions of backward and forward propagating electromagnetic waves depending on the eigenvalue of $U(\Lambda_{\infty})$ and this dependency gives rise to correlations between the values of local observables identical to those experienced with an entangled state of two qubits. We then describe an experiment capable of distinguishing the two eigenvalues of $U(\Lambda_{\infty})$ providing experimental falsification of the theory. Using the above experiment as a primitive we introduce a model of quantum computation in which the logical qubit is not postulated as a two level system, but arises as the representation space of quantum states describing the entanglement between degrees of freedom of a single photon (e.g. helicity and direction of propagation). We call the resource of the computation single photon entanglement qubit and show it is canonically isomorphic to a standard qubit: its pure states form a Bloch sphere, its logical observables close an $\mathfrak{su}(2)$ algebra, and its admissible physical transformations act as automorphisms of the Bloch sphere. We then construct single photon entanglement qubits entangling gate exploiting parity measurements of physical degrees of freedom. Finally, we establish universality of the computational model in the conventional circuit sense.
@@ 1 Introduction
Quantum entanglement is usually introduced as a feature of composite quantum systems: given a tensor product $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, there exist pure states that cannot be written as product vectors. In nonrelativistic quantum mechanics this structure is taken as primitive. By contrast, in relativistic quantum theory the state space is constrained by spacetime symmetry: one starts from unitary representations of the Poincaré group and builds quantum fields and particles from them. It is then natural to ask whether some instances of entanglement may have a direct geometric origin in the representation theory of spacetime symmetries.
@@
$\mathcal{L}_{\mathrm{ext}}\;\cong\;\bigl(\mathrm{SO}(3,1)^{+}_{\uparrow}\rtimes_{\mathrm{Ad}_{\Lambda_{\infty}}}\mathbb{Z}_{2}\bigr)\times\mathbb{Z}_{2}$
shares the same identity component as the ordinary Lorentz group $O(3,1)$ but has disconnected components generated by $\Lambda_{\infty}(\theta,\phi)$ and $\Lambda_{-\infty}(\theta,\phi)$ in place of parity and time-reversal. When translations are included, this leads to an extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{1,3}\rtimes\mathcal{L}_{\mathrm{ext}}$ whose unitary irreducible representations (UIRs) differ from the standard Wigner classification [2] in two essential ways: (i) timelike and spacelike orbits are merged into a single tachyon/massive multiplet, and (ii) in the massless sector the stability subgroup is enlarged from Wigner’s $ISO(2)$ to
$ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}\mathbb{Z}_{2},\qquad\mathbb{Z}_{2}=\{1,-1\},\quad\Lambda_{-\infty}^{2}=I.$
The presence of this extra $\mathbb{Z}_{2}$ factor in the stabiliser of a lightlike orbit of the extended group has a direct and unavoidable representation–theoretic consequence. Every massless UIR of $\mathcal{P}_{\mathrm{ext}}$ is no longer a single forward Wigner representation, but comes as a two components object
@@
where $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are forward and backward lightlike Wigner’s UIRs of ordinary Poincaré group, and $\varepsilon$ is the eigenvalue of the unitary operator $U(\Lambda_{\infty})$. Note that $\Lambda_{-\infty}$ is the element in the little group that fixes $p_{0}$ enlarging the stability group of lightlike orbit from $ISO(2)$ to $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}$, while $\Lambda_{\infty}$ exchanges the forward and backward sectors; on $\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}}$ the latter is represented by an off-diagonal involution whose eigenvectors have eigenvalues $\varepsilon=\pm 1$. In consequence of this the massless sector always carries an additional binary internal label $\varepsilon=\pm 1$ attached to the pair $(\pi^{+},\pi^{-})$.
In section 2.3 we show that this binary structure realises, in a spacetime–geometric way, the algebraic pattern usually associated with entangled qubit pairs. The representation space of $\pi_{\mathrm{ml}}^{\mathrm{ext}}$ can be written as a direct sum
@@
where the above summands carry forward and backward massless Wigner UIRs respectively, while the action of $U(\Lambda_{\infty})$ exchanges the two sectors up to a sign determined by $\varepsilon$. In particular we show that there exist:
@@
In section 3 we show that these massless representations have already been realized in laboratory exploiting entanglement in degrees of freedom of a single photon (see e.g. [4]).
We then explore these photonic states as a resource for quantum computation. Quantum computation is usually presented by postulating a two–level system as a primitive. In photonics this is often realised by selecting a two–dimensional code subspace inside a larger physical Hilbert space (e.g. polarisation within a multi–mode field, dual–rail in spatial modes, time bins, etc.). In section 4 we show that the representation space of quantum states describing the entanglement between degrees of freedom of a single photon (e.g. helicity and direction of propagation) already identify the quantum states of a logical qubit. Operationally, the resulting qubit is implemented as a single photon entanglement (SPE) state between propagation direction and polarisation and the two logical basis states appear as Bell–like states in a fixed two–mode single–photon manifold. The qubit code space $\mathcal{C}$ is operatively constructed by providing the single photon entanglement logical basis with two physically standard one–parameter optical controls (a relative phase shifter and a coherent two–mode coupler) that, when represented in the logical basis, generate all single–qubit unitaries $SU(2)$ on the code space $\mathcal{C}$. This choice immediately provides a canonical Bloch sphere of pure logical states and a CPTP instrument on the tensor product space of two of such Bloch spheres whose Choi state provides an entangling resource for universal quantum computation. In particular, in section 5, we will introduce a two–photons polarisation parity measurement and show that expressed in logical basis is equivalent to a Pauli–product measurement on the tensor product of two single photon entanglement (SPE) qubits $\mathcal{C}_{A}\otimes\mathcal{C}_{B}$, from which an entangling logical gate follows via a standard gate–teleportation construction. From this it follows universality of the computational model.
2 Massless Representations of Extended Poincarè Group
2.1 Unitary Irreducible Representations (UIRs) Classification
@@
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. Including translations we defined the extended Poincaré group
@@
on UIRs of Poincaré group $\mathcal{P}_{0}$. In order to explain this we call $N\equiv\mathcal{P}_{0}=(\mathcal{T}\rtimes SO(3,1)^{+}_{\uparrow})$ and define
@@
An element $z\in Z$ acts on $N$ via automorphisms, namely its action on a given element of the Poincaré group $n=(h,a)\in N$ with $h\in SO(3,1)^{+}_{\uparrow}$:
@@
This induces an action on the set of characters (i.e. UIRs) of $\mathcal{P}_{0}$:
@@
The set $\hat{N}$ is constituted by UIRs of Poincaré group which have been classified by Wigner in [2]. UIRs of Poincaré group are classified by the value of the four momentum vector $p=p_{\mu}$ together with the value of the invariant lenght $|L_{s}\cdot p|^{2}$ associated to the orbit
@@
$p$ is called the representative of the induced UIR of Poincaré group. Given choice of $p_{\mu}$ an $SO(3,1)^{+}_{\uparrow}$-orbit representative in Wigner’s classification, its invariant could be $p^{\mu}p_{\mu}\neq 0$ or $p^{\mu}p_{\mu}=0$. Choosing a non massless ($p^{\mu}p_{\mu}\neq 0$) representative $p_{\mu}$ we have:
@@
2.2 Massless UIRs
Derivation of massless UIRs of the extended group are detailed in [1]. Here we summarize the relevant conceptual pillars. It is considered the lightlike zz–orbit
@@ 2.2 Massless UIRs
thus interchanging the forward and backward light–cones. At the level of massless UIRs of $\mathcal{P}_{\text{ext}}$, $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent since the unitary operator representing $U(\Lambda_{\infty})$ is an intertwiner between $\pi_{\text{fwd}}$ and $\pi_{\text{bwd}}$
@@
Thus, under the action of the discrete factor $Z_{(\theta,\phi)}=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ on the dual space $\widehat{\mathcal{P}_{0}}$, (i.e. the set of Wigner’s UIRs) the set $\Pi\equiv\{\pi^{\text{fwd}},\pi^{\text{bwd}}\}$ constitute a single equivalence class of UIRs in which the transformation $\Lambda_{\infty}$ is a non trivial stabilizer, namely:
@@
where $C$ implements the intertwining between $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ as above and its representation depends on the massless representation choosen for Wigner’s little group $ISO(2)$. Throughout this paper we will use helicity representation as working hypothesis and set $C=I_{\mathcal{H}}$ choosing the identification $\mathcal{H}\simeq\mathcal{H}_{\text{bwd}}\simeq\ H_{\text{fwd}}$. Using a section $k_{0}$ adapted as
@@
thus $\bar{U}(\Lambda_{\infty})$ correctly represents the extension.
@@
2.3 Operational Equivalence of Massless UIRs with Entangled Two–Qubits States
@@
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=(I\otimes\sigma_{x})\big(\psi_{\text{fwd}}\otimes|\text{fwd}\rangle+\psi_{\text{bwd}}\otimes|\text{bwd}\rangle\big)$ (39)
@@
Since this must hold for all wavefunctions in $\mathcal{H}_{\oplus}$ we conclude that, via the isometric identification $V$, the sector–swap operator $U(\Lambda_{\infty})$ on $\mathcal{H}\oplus\mathcal{H}$ is unitarily equivalent to the operator $I\otimes\sigma_{x}$ on $\mathcal{H}\otimes\mathbb{C}^{2}$.
3 Experimental Access to Massless UIRs of Extended Poincarè Group
@@ 3.1 Specialisation to a Fixed Direction and Single–Mode Photons
which are the forward and backward representatives associated with this choice of $\hat{n}$. Rotating the entire setup in space would simply correspond to choosing a different pair $(\theta,\phi)$ in the abstract construction. So $p_{0}=\omega(1,0,0,1),$ with $\Lambda_{\infty}=\Lambda_{\infty}(0,0)$ an involution exchanging the time coordinate $t$ with the spatial coordinate $z$. we can represent a fixed helicity component by kets